# Unfolding AlphaFold’s Bayesian Roots in Probability Kinematics

### v2

@@ Abstract
We present a novel theoretical interpretation of AlphaFold1 that reveals the potential of generalized Bayesian updating for probabilistic deep learning. The seminal breakthrough of AlphaFold1 in protein structure prediction by deep learning relied on a learned potential energy function, in contrast to the later end-to-end architectures of AlphaFold2 and AlphaFold3. While this potential was originally justified by referring to physical potentials of mean force (PMFs), we reinterpret AlphaFold1’s potential as an instance of probability kinematics — also known as Jeffrey conditioning — a principled but under-recognised generalization of conventional Bayesian updating. Probability kinematics accommodates uncertain or soft evidence in the form of updated probabilities over a partition. This perspective reveals AlphaFold1’s potential as a form of generalized Bayesian updating, rather than a thermodynamic potential. To confirm our probabilistic framework’s scope and precision, we analyze a synthetic 2D model in which an angular random walk prior is updated with evidence on distances via probability kinematics, mirroring AlphaFold1’s approach. This theoretical contribution connects AlphaFold1 to a broader class of well-justified Bayesian methods, allowing precise quantification, surpassing merely qualitative heuristics based on PMFs. Our contribution is theoretical: we replace AlphaFold1’s heuristic analogy with a principled probabilistic framework, tested in a controlled synthetic setting where correctness can be assessed. More broadly, our results point to the considerable promise of probability kinematics for probabilistic deep learning, by allowing the formulation of complex models from a few simpler components.
@@ 1 Introduction
The protein structure prediction problem – predicting the three-dimensional (3D) structure of a protein from its amino acid sequence – has been a central unsolved challenge in molecular biology for more than 50 years [17]. In 2024, the shared Nobel Prize in Chemistry was awarded to John Jumper and Demis Hassabis for AlphaFold’s solution of this problem [10]. Here, we focus on the initial formulation of AlphaFold, which we will refer to as AlphaFold1 (AF1) [55, 56]. AF1 differs considerably from the end-to-end prediction approach of its two successors [34, 1] in that it makes use of a potential function parameterized by deep learning. Predictions by AF1 are obtained from minimizing this potential energy function. We propose here that AF1’s general approach is well worth reconsidering as a basis for further development, as we discuss below.
AF1’s potential function was justified by qualitative analogy to physical potentials of mean force (PMFs) [55, 56] as used in the physics of liquids, first formulated by Kirkwood in 1935 [37]. Here, we show instead that AF1 is an instance of an under-recognized generalization of Bayesian updating called Jeffrey conditioning or probability kinematics (PK). PK was introduced by the philosopher of logic and probability Richard C. Jeffrey in 1957 [31, 32, 33], and specifies how to incorporate soft evidence in the form of updated probabilities on a partition of a random variable (see Figure 1), rather than a concrete event. AF1 updates an empirical Bayes prior over protein dihedral angles with soft evidence on pairwise distances between amino acids, both obtained from deep models.
$\xi_{1}$$\xi_{2}$$\xi_{3}$$\xi_{4}$$\xi_{5}$$\xi_{6}$$\xi_{7}$
@@
  • Building on a definition of PK for PDFs and infinite partitions tailored to the case of AF1, we propose a generalized Bayesian reinterpretation of AF1’s potential function.
  • We contrast our framework with AF1’s heuristic explanation based on PMFs and link its deep model to statistical models of protein structure. This bridge is timely, as there is a growing need for modelling epistemic and aleatoric uncertainty [9, 53, 11, 2, 21, 12, 42].
  • We introduce a tractable two-dimensional synthetic model that is analogous to AF1’s potential. The model updates an angular prior with evidence in the form of a PDF over a distance. This setup mirrors the probabilistic framework behind AF1’s potential and allows us to demonstrate the update’s precision via hypothesis testing.
  • We reveal PK as a previously unrecognised foundation for formulating deep generative models of biomolecular structure.
  • Our results show PK’s relevance and potential in formulating compositional deep generative models, ie. models constructed by combining simpler probabilistic components into a coherent whole [22, 69, 6].
Our aim is not to improve or asses empirical accuracy, as AF1’s historical success is well established [55, 56], but to clarify its previously unrecognized statistical foundation. To do this, we replace AF1’s heuristic analogy with a principled probabilistic framework, tested in a controlled synthetic setting where correctness can be assessed.
@@ 2.1 Preliminaries
Figure 2: (left) Two amino acids in a protein, with relevant atoms and dihedral angles labeled. Atoms that do not play a role in the AF1 potential, including side chain and backbone oxygen atoms, are shown in white. The continuation of the protein backbone is indicated with dots. (right) A protein (Protein Data Bank identifier 1L2Y) with a subset of all the $C\beta-C\beta$ distances shown in black dotted lines. Only the atoms relevant to the AF1 potential are shown. Figures made with PyMol [54].
Notation and assumptions
We will use the following generic notation to emphasize features common to PK and PMFs. Let $\boldsymbol{\omega}$ be a random variable, and let $\boldsymbol{\xi}$ be a random variable obtained via a many-to-one mapping, $\boldsymbol{\xi}(\boldsymbol{\omega})$. Thus, $\boldsymbol{\xi}$ is a random variable and $\boldsymbol{\xi}(\boldsymbol{\omega})$ is the specific realization of $\boldsymbol{\xi}$ for a given value of $\boldsymbol{\omega}$. We will call $\boldsymbol{\omega}$ a fine-grained random variable, and $\boldsymbol{\xi}$ a corresponding coarse-grained or collective random variable. Depending on the context, $\boldsymbol{\omega}$ and $\boldsymbol{\xi}$ may be vector-valued, scalar or discrete; in the latter two cases we omit the boldface for clarity. These symbols are reused and explained in their context as needed.
@@
The protein amino acid sequence, $\boldsymbol{s}$, of length $L,$ is represented by a $20L$ vector of concatenated 20-dimensional one hot encodings of the 20 amino acid symbols. $\boldsymbol{M}$ is an $M\times 20L$ matrix that contains the multiple sequence alignment of $M$ protein sequences related to $\boldsymbol{s}$. A protein structure will be represented by its backbone atoms only, including $N,C\alpha,C,C\beta$ atoms (see Figure 2). The 3D coordinates of these four atoms for each amino acid are contained in the tensor, $\boldsymbol{X},$which is $L\times 4\times 3$.
@@
Protein structure is often parameterized using dihedral angles, [^1] A dihedral angle or torsion angle, defined by four consecutively bonded atoms, is the angle between the planes formed by the first three atoms and the last three. Under the common assumption that bond lengths and bond angles take fixed, ideal values, a protein’s 3D structure can be fully parameterized by its set of backbone and side-chain dihedral angles. which has the advantage of invariance to rotation and translation and statistical simplicity [26]. Let $\boldsymbol{\phi}$ and $\boldsymbol{\psi}$ be vectors of $L$ dihedral angles in $\pi,\pi]$. These vectors establish a rotation-invariant representation of the coordinates – we will write $\boldsymbol{X}(\boldsymbol{\phi},\boldsymbol{\psi})$, where $\boldsymbol{X}$ is determined by the angles up to an arbitrary rotation and translation. Following AF1 [55, 56], we will assume ideal bond angles, ideal bond lengths, the third backbone dihedral angle $\omega=\pi$ for all amino acids, and ignore amino acid side chains and the backbone oxygen atom, without any loss of generality (see Figure 2).
@@
Let the random variable $\boldsymbol{D}$ represent the $L\times L$ distance matrix of distances between $L$ $C\beta$ atoms, with L(L1)/2L(L-1)/2 unique values. Note that although these pairwise distances only involve the $C\beta$ atoms, they are taken as representative distances between amino acids. A given set of atomic coordinates, $\boldsymbol{X}$, or dihedral angles, $(\boldsymbol{\phi},\boldsymbol{\psi})$, determines $\boldsymbol{D}$ fully – we will write $\boldsymbol{D}(\boldsymbol{X})$ and $\boldsymbol{D}(\boldsymbol{\phi},\boldsymbol{\psi})$ for the concrete realization of $\boldsymbol{D}$ for given coordinates or dihedral angles. NN will denote a neural network.
@@ 2.2 AF1’s potential
AF1’s potential contains four terms, three of which are parameterized by deep models. A fourth, repulsive potential is derived from physical considerations and merely serves to prevent atoms from overlapping. The first three potentials are central to this article. The exact form of the PDFs mentioned below will be detailed in Section 4.
@@
$V_{\textrm{Dihedral}}(\boldsymbol{\phi},\boldsymbol{\psi})=-\sum_{i}\log p\left(\phi_{i},\psi_{i}\mid\boldsymbol{s},\boldsymbol{M}\right).$ (1)
where the sum runs over all amino acids. The second energy terms contains the pairwise distances between the $C\beta$ atoms, which serve as representative distances between amino acids $i$ and $j$,
$V_{\textrm{Distance}}(\boldsymbol{D})=-\sum_{i,j,i\neq j}\log p\left(D_{i,j}\mid\boldsymbol{s},\boldsymbol{M}\right).$ (2)
As we will show in Section 4, the formulation of a consistent joint probability distribution over angles and distances requires a third term – a so-called reference potential, which also concerns pairwise distances. This potential depends only on the length of the sequence, $L(\boldsymbol{s})$, and a one-hot vector that specifies the positions of Glycine residues, $\mathbb{I}_{\textrm{Glycine}}(\boldsymbol{s})$. Glycine is a small amino acid that allows exceptional conformational freedom to the protein chain. The reference potential is
$V_{\textrm{Reference}}(\boldsymbol{D})=-\sum_{i,j,i\neq j}\log p\left(D_{i,j}\mid L(\boldsymbol{s}),\mathbb{I}_{\textrm{Glycine}}(\boldsymbol{s})\right).$ (3)
What AF1 refers to as a PMF is then obtained from their difference,
$V_{\textrm{PMF}}\left(\boldsymbol{D}\right)\text{=}V_{\textrm{Distance}}\left(\boldsymbol{D}\right)-V_{\textrm{Reference}}\left(\boldsymbol{D}\right).$ (4)
The final potential brings the three potential above together, and combines it with the predefined repulsive potential,
$V_{\textrm{AF1}}(\boldsymbol{\phi},\boldsymbol{\psi})=V_{\textrm{PMF}}\left(\boldsymbol{D}(\boldsymbol{\phi},\boldsymbol{\psi})\right)+V_{\textrm{Dihedral}}(\boldsymbol{\phi},\boldsymbol{\psi})+V_{\textrm{Repulsive}}\left(\boldsymbol{X}(\boldsymbol{\phi},\boldsymbol{\psi})\right).$ (5)
Note that $\boldsymbol{D}$ and $\boldsymbol{X}$ (up to an arbitrary rotation) in the above expression are fully determined by the dihedral angles, $\boldsymbol{\phi}$ and $\boldsymbol{\psi}$ (see Figure 2). Prediction by AF1 amounts to minimizing the potential as a function of the dihedral angles using gradient descent. In the next section, we explain how AF1 justifies this potential by heuristic analogy with PMFs for liquid systems.
@@ 2.3 AF1’s potential justified as a PMF
A PMF can be understood as an "averaged", “mean” or "effective" potential regarding some coordinates (typically called the collective or coarse-grained variables) of a physical system, while averaging over the remaining degrees of freedom [14, 72, 28]. A PMF is defined as
$p(\boldsymbol{\xi})$ $\propto\exp(-\textrm{PMF}(\boldsymbol{\xi})/kT)=\int\delta\left(\boldsymbol{\xi}-\boldsymbol{\xi}(\boldsymbol{X})\right)\exp(-U(\boldsymbol{X})/kT)\textrm{d}\boldsymbol{X},$ (6)
where $k$ is Boltzmann’s constant, $T$ is the temperature, $\boldsymbol{X}$ are the atomic coordinates of the system, $\boldsymbol{\xi}$ is the collective random variable, $\boldsymbol{\xi}(\boldsymbol{X})$ is the realization of $\boldsymbol{\xi}$ at $\boldsymbol{X}$, $\delta(.)$ is the Dirac delta function, and the exponential factor is the Boltzmann distribution associated with internal energy $U.$
PMFs were introduced by Kirkwood in his seminal paper from 1935 [37], applied to mixtures of liquids. What is relevant here is that Kirkwood potentials give rise to expressions of the form,
$\textrm{PMF}_{\textrm{K}}(r)=-kT\log\left(\frac{p(r)}{p_{\textrm{R}}(r)}\right),$ (7)
where $r$ is the distance between two liquid particles. The numerator in the log is the probability density observed in a liquid, while the denominator refers to a so-called reference distribution. In the case of Kirkwood PMFs, the reference distribution is well-defined, corresponding to an ideal state of non-interacting particles [37]. We now turn to the heuristic adoption of formulas similar to Kirkwood PMFs, as done by AF1, for biomolecular prediction and simulation.
@@
Knowledge based potentials (KBPs) attempt to formulate empirical energy functions based on statistics derived from the database of known structures [52, 64, 49, 58]. KBPs based on the pairwise distances between $C\beta$ atoms were introduced in Sippl (1990) [58] and widely adopted [61, 59, 62]. The energy of a protein structure following such a KBP is given by
@@
These potentials were formulated by qualitative analogy with the Kirkwood PMFs discussed in the previous section. Despite undisputed success, their theoretical justification as PMFs was widely contested [65, 5, 38]. Notably, unlike physical PMFs, KBPs are not derived from a Boltzmann distribution of a single system, but from a heterogeneous data set of different protein structures, and the reference distribution is undefined [5, 38]. Attempts at qualitative justification invoked properties of random polymers [20] and probabilistic reasoning [57]. Subsequently, pairwise KBPs were quantitatively explained [25, 67], and eventually recognised as an instance of PK [27], to which we turn next.

### v3

# AlphaFold’s Bayesian Roots in Probability Kinematics

@@ Abstract
The seminal breakthrough of AlphaFold in protein structure prediction relied on a learned potential energy function parameterized by deep models, in contrast to its successors AlphaFold2 and AlphaFold3, which lack an explicit probabilistic interpretation. While AlphaFold’s potential was originally justified by heuristic analogy to physical potentials of mean force, we show that it can instead be understood as a principled instance of probability kinematics (PK), also known as Jeffrey conditioning, a generalization of Bayesian updating. This reinterpretation reveals that AlphaFold is a generalized Bayesian model that explicitly defines a posterior distribution over structures, providing a deeper explanation of its success and a foundation for future model design. To demonstrate this framework with precision, we introduce a tractable synthetic model in which an angular random walk prior is updated with distance-based evidence via PK, directly mirroring AlphaFold’s mechanism. This setting allows us to explore the probabilistic foundations of AlphaFold in a clear and interpretable way. Our work connects a landmark in protein structure prediction to a broader class of compositional deep generative models and points to new opportunities for principled probabilistic approaches.
@@ 1 INTRODUCTION
The protein structure prediction problem – predicting the three-dimensional (3D) structure of a protein from its amino acid sequence – has been a central unsolved challenge for more than 50 years (Dill and MacCallum, 2012). In 2024, the shared Nobel Prize in Chemistry was awarded to John Jumper and Demis Hassabis for AlphaFold’s solution of this problem (Callaway, 2024). Here, we focus on the initial formulation of AlphaFold, which we will refer to as AlphaFold1 (AF1) (Senior et al., 2019, 2020). AF1 makes use of a potential function parameterized by deep learning and (as we show here) rooted in generalized Bayesian updating, which differs considerably from the end-to-end deep learning pipeline approach of its two successors that despite increased pragmatic accuracy lack an explicit, principled, SE(3)-invariant probabilistic model over the conformational space of proteins (Jumper et al., 2021; Abramson et al., 2024; Jing et al., 2025). Predictions by AF1 are obtained by minimizing this potential energy function with respect to the protein’s dihedral angles. We argue that AF1’s probabilistic nature is well worth reconsidering as a basis for further development, as we discuss below.
AF1’s potential function was justified heuristically by appealing to physical potentials of mean force (PMFs) (Senior et al., 2019, 2020) as used in the physics of liquids, first formulated by Kirkwood in 1935 (Kirkwood, 1935). Here, we instead interpret AF1 as an instance of an under-recognized generalization of Bayesian updating called Jeffrey conditioning or probability kinematics (PK). PK was introduced by the philosopher of logic and probability Richard C. Jeffrey in 1957 (Jeffrey, 1957, 1990, 2004), and specifies how to incorporate soft evidence, i.e., updated probabilities on a partition of a random variable’s sample space (see Figure 1), rather than conditioning on a concrete event. AF1 updates a prior distribution over a protein’s dihedral angles with soft evidence on pairwise distances between amino acids, both obtained from deep models.
$\xi_{1}$$\xi_{2}$$\xi_{3}$$\xi_{4}$$\xi_{5}$$\xi_{6}$$\xi_{6}$
@@
  • Building on a definition of PK for probability density functions (PDFs) and infinite partitions tailored to the case of AF1, we propose a generalized Bayesian reinterpretation of AF1´s potential function, and make explicit its functional form as a prior density adjusted by a ratio of PDFs.
  • We contrast our framework with AF1´s heuristic explanation based on PMFs and link its deep model to other statistical models of protein structure. This bridge is timely, as there is a growing need for modeling epistemic and aleatoric uncertainty (Bowman, 2024; Ramasamy et al., 2025; Chakravarty et al., 2024; Agarwal and McShan, 2024; Gavalda-Garcia et al., 2025; Chakravarty et al., 2025; López-Sagaseta and Urdiciain, 2025).
  • We introduce a tractable two-dimensional synthetic model that is analogous to AF1’s potential. The model updates an angular prior with evidence in the form of a PDF over a distance. This setup mirrors the probabilistic framework behind AF1’s potential and allows us to demonstrate the update’s precision via hypothesis testing. In the Appendix, we also provide an application to a small 3D protein model.
  • We reveal PK as a previously unrecognized foundation for formulating deep generative models of biomolecular structure.
  • Our results show PK’s relevance and potential in formulating compositional deep generative models, i.e., models constructed by combining simpler probabilistic components into a coherent whole (Ghahramani, 2015; Wood et al., 2014; Bingham et al., 2019).
Our aim is not to improve or assess empirical accuracy, as AF1’s historical success is well established (Senior et al., 2019, 2020), but to clarify its previously unrecognized statistical foundation and point out its potential for future work. To do this, we make use of a principled probabilistic framework, tested in a controlled synthetic setting where accuracy can be assessed.
@@ 2.1 Preliminaries
Figure 2: (top) Two amino acids in a protein, with relevant atoms and dihedral angles labeled. Atoms that do not play a role in the AF1 potential, including side chain and backbone oxygen atoms, are shown in white. The continuation of the protein backbone is indicated with dots. (bottom) A protein with some $C\beta-C\beta$ distances shown as dotted lines. Only the atoms relevant to the AF1 potential are shown. Figures made with PyMol (Schrödinger, LLC, 2015).
@@
The protein amino acid sequence, $\boldsymbol{s}$, of length $L,$ is represented by a $L\times 20$ tensor of one-hot encodings of the 20 amino acid symbols. $\boldsymbol{M}$ is an $M\times L\times 20$ tensor that contains the multiple sequence alignment of $M$ protein sequences related to $\boldsymbol{s}$. Following AF1, a protein structure will be represented by its backbone atoms only, including $N,C\alpha,C,C\beta$ atoms (see Figure 2). The 3D coordinates of these four atoms for each amino acid are in the tensor $\boldsymbol{X}$, which is $L\times 4\times 3$.
@@
Protein structure is often parameterized using dihedral angles (Ramachandran et al., 1963), [^1] A dihedral angle or torsion angle, defined by four consecutively bonded atoms, is the angle between the planes formed by the first three atoms and the last three. Under the common assumption that bond lengths and bond angles take fixed, ideal values, a protein’s 3D structure is fully determined by its dihedral angles. which has the advantage of invariance to rotation and translation and statistical simplicity (Edgoose et al., 1998; Boomsma et al., 2008; Hamelryck et al., 2012). Let $\boldsymbol{\phi}$ and $\boldsymbol{\psi}$ be vectors of $L$ dihedral angles in $(-\pi,\pi]$. We will write $\boldsymbol{X}(\boldsymbol{\phi},\boldsymbol{\psi})$ for the realization of $\boldsymbol{X}$ up to an arbitrary rotation and translation for a given set of dihedral angles. Following AF1 (Senior et al., 2019, 2020), we assume ideal bond angles and lengths, trans peptide bonds, and ignore amino acid side chains and backbone oxygen atoms, without any loss of generality (see Figure 2).
@@
Let $\boldsymbol{D}$ represent the $L\times L$ distance matrix of distances between $L$ $C\beta$ atoms, with L(L1)/2L(L-1)/2 unique values. Note that although these pairwise distances only involve the $C\beta$ atoms, they are taken as representative distances between amino acids. A given set of atomic coordinates, $\boldsymbol{X}$, or dihedral angles, $(\boldsymbol{\phi},\boldsymbol{\psi})$, determines $\boldsymbol{D}$ fully – we will write $\boldsymbol{D}(\boldsymbol{X})$ and $\boldsymbol{D}(\boldsymbol{\phi},\boldsymbol{\psi})$ for the realization of $\boldsymbol{D}$ for given coordinates or dihedral angles.
We will use the following generic notation to emphasize features common to PK and PMFs. Let $\boldsymbol{\omega}$ be a random variable, and let $\boldsymbol{\xi}$ be a random variable obtained via a many-to-one mapping of $\boldsymbol{\omega}$. Thus, $\boldsymbol{\xi}$ is the random variable and $\boldsymbol{\xi}(\boldsymbol{\omega})$ is its concrete realization for a given value of $\boldsymbol{\omega}$. We will call $\boldsymbol{\omega}$ a fine-grained random variable, and $\boldsymbol{\xi}$ a corresponding coarse-grained or collective random variable. Depending on the context, $\boldsymbol{\omega}$ and $\boldsymbol{\xi}$ may be matrices, vectors, sets or scalars; in the latter two cases we omit the boldface. These symbols are reused and explained in their context as needed. Notably, in the previous paragraph, $\boldsymbol{D}$ is a coarse grained, collective variable, while $\boldsymbol{X}$ is a corresponding fine-grained variable, corresponding to instances of $\boldsymbol{\xi}$ and $\boldsymbol{\omega}$, respectively. Appendix (A) gives an overview of notation and abbreviations.
@@ 2.2 AF1’s potential
AF1’s potential contains four terms, three of which are parameterized by deep models. A fourth, repulsive potential is derived from physical considerations and merely serves to prevent atoms from overlapping. The first three potentials are central to this article. Section 4 provides details on the PDFs mentioned below.
@@
$V_{\textrm{A}}(\boldsymbol{\phi},\boldsymbol{\psi})=-\sum_{i}\log p\left(\phi_{i},\psi_{i}\mid\boldsymbol{s},\boldsymbol{M}\right),$ (1)
where the sum runs over all amino acids. The second energy term contains the pairwise distances between the $C\beta$ atoms, which serve as representative distances between amino acids $i$ and $j$,
$V_{\textrm{D}}(\boldsymbol{D})=-\sum_{i,j,i\neq j}\log p\left(D_{i,j}\mid\boldsymbol{s},\boldsymbol{M}\right).$ (2)
As we will show in Section 4, the formulation of a consistent joint probability distribution over angles and distances requires a third term – a so-called reference potential, which also concerns pairwise distances. In AF1, this potential depends only on the number of amino acids in the protein, $L$, and a one-hot vector that specifies the positions of Glycine residues, $\mathbb{I}_{\textrm{Gly}}(\boldsymbol{s})$. Glycine is a small amino acid that allows exceptional conformational freedom to the protein chain. The reference potential is
$V_{\textrm{R}}(\boldsymbol{D})=-\sum_{i,j,i\neq j}\log p\left(D_{i,j}\mid L,\mathbb{I}_{\textrm{Gly}}(\boldsymbol{s})\right).$ (3)
What AF1 refers to as a Kirkwood PMF is then obtained from their difference,
$V_{\textrm{PMF}}\left(\boldsymbol{D}\right)\text{=}V_{\textrm{D}}\left(\boldsymbol{D}\right)-V_{\textrm{R}}\left(\boldsymbol{D}\right).$ (4)
The final potential brings the three potentials above together, and combines it with the predefined repulsive potential,
$V(\boldsymbol{\phi},\boldsymbol{\psi})=V_{\textrm{PMF}}\left(\boldsymbol{D}(\boldsymbol{\phi},\boldsymbol{\psi})\right)+V_{\textrm{A}}(\boldsymbol{\phi},\boldsymbol{\psi})\\ +V_{\textrm{Repulsive}}\left(\boldsymbol{X}(\boldsymbol{\phi},\boldsymbol{\psi})\right).$ (5)
Note that $\boldsymbol{D}$ and $\boldsymbol{X}$ (up to an arbitrary rotation and translation) in the above expression are fully determined by the dihedral angles, $\boldsymbol{\phi}$ and $\boldsymbol{\psi}$ (see Figure 2). Prediction by AF1 amounts to minimizing the potential as a function of the dihedral angles using gradient descent. In the next section, we explain how AF1 justifies this potential by heuristic analogy with PMFs for liquid systems.
@@ 2.3 AF1’s potential justified as a PMF
A PMF can be understood as an averaged, mean or effective potential regarding some coordinates (typically called the collective or coarse-grained variables) of a physical system, while averaging over the remaining degrees of freedom (Chandler, 1987; Zuckerman, 2010; Hartmann et al., 2011). A PMF is defined as
$p(\boldsymbol{\xi})$ $\propto\exp(-\textrm{PMF}(\boldsymbol{\xi})/kT)$ (6)
$=\int\delta\left(\boldsymbol{\xi}-\boldsymbol{\xi}(\boldsymbol{X})\right)\exp(-U(\boldsymbol{X})/kT)\textrm{d}\boldsymbol{X},$ (7)
where $k$ is Boltzmann’s constant, $T$ is the temperature, $\boldsymbol{X}$ are the atomic coordinates of the system, $\boldsymbol{\xi}$ is the collective random variable (e.g., distances between a subset of the atoms), $\boldsymbol{\xi}(\boldsymbol{X})$ is the realization of $\boldsymbol{\xi}$ at $\boldsymbol{X}$, $\delta(.)$ is the Dirac delta function, and the exponential factor is the Boltzmann distribution associated with the internal energy $U$.
PMFs were introduced by Kirkwood in his seminal paper from 1935 (Kirkwood, 1935), applied to mixtures of liquids. For this particular physical system, Kirkwood proposed expressions of the form,
$W_{\textrm{Kirkwood}}(r)=-kT\log\left(\frac{p(r)}{p_{\textrm{Reference}}(r)}\right),$ (8)
where $r$ is the distance between two identified liquid particles. The numerator in the log is the probability density observed in a liquid, while the denominator refers to a so-called reference distribution. In the case of Kirkwood potentials, the reference distribution is well-defined, corresponding to an ideal state of non-interacting particles (Kirkwood, 1935). It serves to isolate genuine intermolecular interactions from features that are purely geometric or entropic. Technically, a Kirkwood potential is thus a relative PMF, corresponding to a difference between the PMFs of the interacting liquid and its corresponding non-interacting reference system. We now turn to the heuristic adoption of formulas resembling Kirkwood potentials, as done by AF1, for biomolecular prediction and simulation.
@@
Knowledge based potentials (KBPs) attempt to formulate empirical energy functions based on statistics derived from the database of known structures (Pohl, 1971; Tanaka and Scheraga, 1976; Miyazawa and Jernigan, 1985; Sippl, 1990). KBPs based on the pairwise distances between $C\beta$ atoms were introduced by Sippl (1990) and widely adopted (Sippl, 1995, 1993a; Sippl et al., 1996). The energy of a protein structure following such a KBP is given by
@@
where $s_{i},s_{j}$ are amino acid labels. These heuristic potentials were loosely inspired by the ratio of PDFs arising in Kirkwood potentials (Eq. 8).
Early Rosetta work (Simons et al., 1997) used a Sippl-style potential involving discretized distances combined with Monte Carlo sampling of protein fragments. Sippl-style potentials were also used for protein structure evaluation, fold recognition and threading, and model refinement (Zhou and Zhou, 2002; Shen and Sali, 2006; Lu et al., 2008; Zhang and Zhang, 2010). In the 2010s, later versions of Rosetta and other programs instead typically relied on predicted contacts/distances as straightforward restraints, which is distinct from Sippl-style potentials and probability kinematics.
Despite pragmatic success, their theoretical justification as genuine PMFs was widely contested (Thomas and Dill, 1996; Ben-Naim, 1997; Koppensteiner and Sippl, 1998). Notably, unlike physical PMFs and Kirkwood potentials, KBPs are not derived from a Boltzmann distribution of a single physical system, but from a heterogeneous data set of different protein structures, and the reference distribution is undefined (Ben-Naim, 1997; Koppensteiner and Sippl, 1998). Attempts at qualitative, heuristic justification invoked properties of random polymers (Finkelstein et al., 1995) and probabilistic reasoning (Simons et al., 1997). Subsequently, KBPs were quantitatively, formally explained (Hamelryck et al., 2010; Valentin et al., 2014), and eventually recognized as an instance of PK (Hamelryck et al., 2015), to which we turn next.